# Images from disk and spherical accretions of Bardeen black hole surrounded by perfect fluid dark matter

### v1

@@ Abstract
In this paper, we investigate the shadow and optical appearance of the Bardeen black hole surrounded by perfect fluid dark matter (PFDM) illuminated by various static accretions. First, we find that as the dark matter parameter $\left|\alpha\right|$ increases, the fundamental characteristic quantities of the black hole, the event horizon $r_{h}$, the photon sphere radius $r_{ph}$, and the critical impact parameter $b_{ph}$ all increase, while the peak of the effective potential $V_{\text{eff}}$ decreases and shifts toward the direction of increasing $r$. In contrast, the magnetic charge parameter $g$ suppresses $r_{h}$, $r_{ph}$, and $b_{ph}$, while increases the peak of $V_{\text{eff}}$ and shifts it toward the direction of decreasing $r$. This indicates a competing effect between the dark matter parameter $\left|\alpha\right|$ and the magnetic charge parameter $g$ on the fundamental properties of the black hole. Furthermore, we use the EHT observational data to constrain the dark matter parameters $\alpha$ and find that the constraint range given by the supermassive black hole SgrA* is stricter than that of the black hole M87*. Finally, the black hole images are studied based on different accretion models, and it is found that both dark matter parameters $\alpha$ and accretion models significantly influence the black hole images. For larger dark matter parameters $\left|\alpha\right|$, the inner shadow or central faint illuminating region of the Bardeen black hole surrounded by PFDM is larger but the bright ring of the image is fainter. This provides a potential method for us to distinguish between classical Bardeen black holes and Bardeen black holes surrounded by PFDM. These preliminary results may provide some clues for future investigations of dark matter using black hole shadows and images.
@@ I Introduction
Recent breakthroughs in Very Long Baseline Interferometry (VLBI) by the Event Horizon Telescope (EHT) collaboration have opened new avenues for probing the fundamental nature of gravity in the strong field regime. The most significant achievement is that EHT successfully captured and published the first images of two supermassive black holes: one is the supermassive black hole at the center of the M87* galaxy EventHorizonTelescope:2019dse; EventHorizonTelescope:2019uob; EventHorizonTelescope:2019jan; EventHorizonTelescope:2019ths; EventHorizonTelescope:2019pgp; EventHorizonTelescope:2019ggy, and the other is the supermassive black hole SgrA* at the center of the Milky Way system EventHorizonTelescope:2022wkp; EventHorizonTelescope:2022apq; EventHorizonTelescope:2022wok; EventHorizonTelescope:2022exc; EventHorizonTelescope:2022urf; EventHorizonTelescope:2022xqj. Due to the strong gravitational lensing effect of the black hole, the surrounding light rays are captured and form the central shadow in the image Virbhadra:1999nm; Virbhadra:2002ju; Virbhadra:2007kw; Synge:1966okc; Bardeen:1972fi; Bozza:2010xqn. The characteristics of the shadow are closely related to the unstable circular orbit of the photon. and the background radiation source. For a parallel light source at infinity, the boundary of the black hole shadow is determined by the critical impact parameter corresponding to the unstable circular orbit of the photon. Perlick:2021aok. Moreover, the bright ring-like structure is the result of the superposition of multiple light rays. This is closely related to the distribution of accretion matter, the structure of accretion disk and strong gravitational lensing effects around the black hole.
@@
In fact, originally, the black hole shadow was termed the escape cone of the photon Synge:1966okc. Subsequently, the angular radius of the photon capture region of the Schwarzschild black hole was derived Luminet:1979nyg. Bardeen’s research further showed that a rotating Kerr black hole would produce a D-shaped shadow structure, which is attributed to the dragging effect of the black hole, resulting in an asymmetry in the orbital radii of prograde and retrograde photons Bardeen:1972fi. This has given rise to extensive discussion on black hole shadows, and various numerical simulations of black hole shadows have been discussed Shen:2005cw; Yumoto:2012kz; Atamurotov:2013sca; Papnoi:2014aaa; Abdujabbarov:2015xqa; Kumar:2018ple, which have shown that black hole shadows depend on spacetime geometry. In addition, black hole shadows in various modified theories of gravity and different spacetime dimensions have been extensively investigated, such as Gauss-Bonnet theory Ma:2019ybz; Guo:2020zmf, conformal gravity Meng:2022kjs, the Chern-Simons type theory Meng:2023wgi; Ayzenberg:2018jip; Amarilla:2010zq, f(R) gravity Addazi:2021pty; Dastan:2016vhb and so on Amarilla:2011fx; Amarilla:2013sj; Amir:2017slq; Mizuno:2018lxz; Eiroa:2017uuq; Vagnozzi:2019apd; Banerjee:2019nnj; Chowdhuri:2020ipb; Konoplya:2019goy; Younsi:2016azx; Olmo:2023lil. Furthermore, the shadows of naked singularities have been discussed Shaikh:2018lcc; Joshi:2020tlq; Dey:2020bgo, and the wormholes shadows were also studied Rahaman:2021web; Kasuya:2021cpk; Shaikh:2018oul; Wielgus:2020uqz; Peng:2021osd; Neto:2022pmu; Tsukamoto:2012xs and references therein. It is worth noting that the EHT has provided observational constraints on the shadow characteristics of the supermassive black holes M87* and SgrA*. For example, EHT observations have constrained the angular diameter of the SgrA* black hole shadow to be $\theta_{d}=48.7\pm 7\mu as$ EventHorizonTelescope:2022wkp, and the angular diameter of the supermassive M87* black hole shadow to be $29.32\mu as<\theta_{d}<51.06\mu as$ Kuang:2022ojj; EventHorizonTelescope:2021dqv. Although the supermassive black holes M87* and SgrA* are generally considered to be Kerr black holes, the limited resolution of the EHT still leaves room for alternative black hole models, or those predicted by modified theories of gravity. In addition, the astronomical observations of the supermassive black holes M87* and SgrA* can be used to test and constrain both alternative general relativistic black hole models and modified theories of gravity. Consequently, using the black hole shadow observed by the EHT to test and constrain the modified gravity theories has been widely discussed in Vagnozzi:2022moj; Afrin:2021imp; Kuang:2022ojj; Tang:2022hsu; Tang:2022bcm; Kuang:2022xjp; Kumar:2019pjp; Shaikh:2021yux; Wu:2023yhp; Capozziello:2023tbo; Sui:2023rfh; Pantig:2022qak; Ghosh:2023kge; Tsukamoto:2014tja.
@@
The bright ring-like structure in EHT images is closely associated with accreting matter around the black hole, and its spatial distribution is significantly influenced by the surrounding luminous accretion flow. This accretion material is mainly composed of gas and dust, forming a rapidly rotating accretion disk around the black hole before ultimately plunging into it. The real astrophysical environment is very complex. The formation of black hole images arises from the combined effects of accretion disk radiation, photon rings, gravitational lensing and relativistic effects, which generally require numerical simulations using general relativistic magnetohydrodynamics (GRMHD) EventHorizonTelescope:2019pcy. In fact, some simple accretion disk models are sufficient to capture key features of black hole images and reveal the properties of strong gravitational field regions. Wald et al. first investigated the imaging features of Schwarzschild black hole using a simplified thin accretion disk model Gralla:2019xty, and defined three types of emission from the accretion disk based on the number of intersections of light rays with the accretion disk: direct emission ($m=1$), lensed ring emission ($m=2$), and photon ring emission ($m\geq 3$). Their study revealed that the observed intensity of the accretion disk primarily originates from direct emission. The contribution of lensed ring emission is relatively small, and the contribution of photon ring emission is negligible. Subsequently, the thin accretion disk models and spherically symmetric accretion models for various types of black holes have been widely studied Dokuchaev:2019pcx; Peng:2020wun; He:2021htq; Eichhorn:2021iwq; Li:2021riw; Wang:2023vcv; Wang:2024lte; Meng:2024puu; Zeng:2020dco; Saurabh:2020zqg; Zeng:2020vsj; Qin:2020xzu; Narayan:2019imo. These studies demonstrate that the black hole images are primarily determined by the properties of spacetime geometry rather than the details of accretion. It is noteworthy that the black holes with double photon sphere structures exhibit additional ring features in their images Gan:2021xdl; Gan:2021pwu; Meng:2023htc. Moreover, considering more realistic astrophysical conditions, the accretion images of various types of rotating black holes have also been extensively studied Hou:2022eev; Wang:2023fge; Liu:2021yev; Zhang:2023bzv; He:2024amh; Zheng:2024ftk; Guo:2024mij; Heydari-Fard:2023kgf; Donmez:2024lfi; Yang:2024nin; Li:2024ctu; Wang:2024uda; Meng:2025ivb. In general, using black hole image features to distinguish between general relativistic black holes Guo:2022iiy; Guo:2022ghl; Archer-Smith:2020hqq; Okyay:2021nnh; Uniyal:2022vdu; Hou:2022eev; Uniyal:2023inx; Akbarieh:2023kjv; Gao:2023mjb; Theodosopoulos:2023ice; Meng:2023uws, other compact objects Boshkayev:2022vlv; Xavier:2023exm; Sakai:2014pga; Bacchini:2021fig; Destounis:2023khj, and black holes in modified gravity theories is a powerful detection method, but this does not always work, as even in GR, the images of black holes may exhibit degeneracies. For example, sufficiently dense boson stars Liebling:2012fv; Rosa:2023qcv may have shadows similar to those of classical black holes Rosa:2022tfv and images of very relativistic rotating boson stars may resemble Kerr black holes Vincent:2015xta.
@@
This paper is organized as follows. In Sec.II, we study the effective potential of photons around the black hole to analyze the photon sphere radius and the impact parameter and also constrain the range of the coupling parameter $\lambda$ by means of EHT observations. In Sec.III, utilizing the Lagrange equation and the ray-tracing method, we calculated the deflection trajectory of photons around the Bardeen black hole surrounded by PFDM. Furthermore, We employed two simplified model functions to investigate the total observed intensity illuminated by the thin disk accretion and the sources of its contribution. In Sec.IV, We presented the shadow and image of a black hole illuminated by a static spherical accretion disk. Sec.V presents the conclusion.
@@ II Photons sphere and observational constraints of Bardeen black hole surrounded by PFDM
For a static spherically symmetric spacetime, considering the coupling between the gravitational field and the non-linear electromagnetic field, a Bardeen black hole surrounded by PFDM can be obtained Zhang:2020mxi
@@
where $g$, $M$ and $\alpha$ are the magnetic charge, mass and dark matter parameters of the black hole respectively. In addition, to ensure the validity of the weak energy condition, the dark matter parameter $\alpha$ must satisfy $\alpha<0$ Sun:2023woa. When the dark matter parameter $\alpha=0$, the black hole reduces to the regular Bardeen black hole. For magnetic charge $g=0$, the black hole reduces to the Schwarzschild black hole surrounded by PFDM Li:2012zx; Ayon-Beato:2000mjt, and when $\alpha=0$, $g=0$, the black hole reduces to the Schwarzschild black hole. In the following study, we will rescale all the physical quantities by the mass $M$ and set $M=1$ for simplicity. By solving the equation $f(r)=0$, we can determine the event horizon radius $r_{h}$ of the black hole. However, since the analytical solution of the event horizon $r_{h}$ is too complicated, we show the relationship between the event horizon $r_{h}$ and the magnetic charge parameter $g$ and the dark matter parameter $\alpha$ in Fig. 1. Obviously, the dark matter parameter $\alpha$ and the magnetic charge parameter $g$ exhibit distinctly different influences on the event horizon. As the coupling parameter $\left|\alpha\right|$ decreases or $g$ increases, the event horizon radius $r_{h}$ shows a monotonically decreasing trend.
Next, we investigate the null geodesic motion and shadow of the Bardeen black hole surrounded by PFDM. The trajectory of photons in curved spacetime can be described by the Euler-Lagrange equation as
@@
$E\equiv-\frac{\partial\mathcal{L}}{\partial\dot{t}}=f(r)\dot{t}~~~~~~~~~~L_{z}=\frac{\partial\mathcal{L}}{\partial\mathcal{\dot{\phi}}}=r^{2}\sin^{2}\theta\dot{\phi}.$ (4)
In addition, we consider the motions of photons moving on the equatorial plane $\theta=\pi/2$ without losing generality. Furthermore, considering $\mathcal{L}=0$ for the photon and defining the impact parameter $b=L_{z}/E$, we can obtain three first-order differential equations of motion for the photons.
@@
where $b_{ph}$ is the critical impact parameter corresponding to the radius $r_{ph}$ of photon sphere. Subsequently, the relationship between the photon sphere radius $r_{ph}$ and the parameters (the magnetic charge $g$ and the dark matter parameter $\alpha$) is shown below in Fig. 1. Similar to the event horizon, the smaller dark matter parameter $\left|\alpha\right|$ or the larger magnetic charge parameter $g$, the smaller the corresponding photon sphere radius $r_{ph}$. In addition, we give the curves of critical impact parameters and effective potential changing with parameters $\alpha$ and $g$ in Fig.2. Obviously, the smaller $\left|\alpha\right|$, the larger $g$, the smaller the critical impact parameter $b_{ph}$ of the black hole, the larger the peak value of the effective potential, and the peak of the effective potential shifts leftward. Overall, for the event horizon $r_{h}$, the radius of the photon sphere $r_{ph}$ and the effective potential $V_{\text{eff}}$, there is a competitive effect between the dark matter parameter $\left|\alpha\right|$ and the magnetic charge parameter $g$.
Figure 1: The event horizon $r_{h}$ and the photon sphere radius $r_{ph}$ as a function of the dark matter parameter $\alpha$ and the magnetic charge parameter $g$ of the Bardeen black hole surrounded by PFDM. Here we set $M=1$.
@@
On the other hand, for the observer at distant positions, the angular diameter $\theta_{d}$ of black hole shadow can be given by Kumar:2020owy
@@
Figure 3: Left: the constraints of the coupling parameters $\alpha$ and $g$ by the observational shadows of M87* black holes. The red curve corresponds to the angular diameter $\theta_{d}=51.06~\mu as$. Right: the constraints of the coupling parameters $\alpha$ and $g$ by the observational shadows of SgrA* black holes. The red curve corresponds to the angular diameter $\theta_{d}=55.7~\mu as$.

### v2

@@ Abstract
In this paper, we investigate the shadow and optical appearance of the Bardeen black hole surrounded by perfect fluid dark matter (PFDM) illuminated by various static accretions. First, we find that as the dark matter parameter $\left|\alpha\right|$ increases, the fundamental characteristic quantities of the black hole, the event horizon $r_{h}$, the photon sphere radius $r_{ph}$, and the critical impact parameter $b_{ph}$ all increase, while the peak of the effective potential $V_{\text{eff}}$ decreases and shifts toward the direction of increasing $r$. In contrast, the magnetic charge parameter $g$ suppresses $r_{h}$, $r_{ph}$, and $b_{ph}$, while increases the peak of $V_{\text{eff}}$ and shifts it toward the direction of decreasing $r$. This indicates a competing effect between the dark matter parameter $\left|\alpha\right|$ and the magnetic charge parameter $g$ on the fundamental properties of the black hole. Furthermore, we use the EHT observational data to constrain the dark matter parameters $\alpha$ and find that the constraint range given by the supermassive black hole SgrA* is stricter than that of the black hole M87*. Finally, we investigate black hole images under different accretion models. The results reveal that both dark matter parameters $\alpha$ and accretion models significantly influence the black hole images. For larger dark matter parameters $\left|\alpha\right|$, the inner shadow or central faint illuminating region of the Bardeen black hole surrounded by PFDM is larger, but the bright ring of the image is fainter. This provides a potential method for us to distinguish between classical Bardeen black holes and Bardeen black holes surrounded by PFDM. These preliminary results may provide some clues for future investigations of dark matter using black hole shadows and images.
@@ I Introduction
Recent breakthroughs in Very Long Baseline Interferometry by the Event Horizon Telescope (EHT) collaboration have opened new avenues for probing the fundamental nature of gravity in the strong field regime. The most significant achievement is that the EHT successfully captured and published the first images of two supermassive black holes: one is the supermassive black hole at the center of the M87* galaxy EventHorizonTelescope:2019dse; EventHorizonTelescope:2019uob; EventHorizonTelescope:2019jan; EventHorizonTelescope:2019ths; EventHorizonTelescope:2019pgp; EventHorizonTelescope:2019ggy, and the other is the supermassive black hole SgrA* at the center of the Milky Way system EventHorizonTelescope:2022wkp; EventHorizonTelescope:2022apq; EventHorizonTelescope:2022wok; EventHorizonTelescope:2022exc; EventHorizonTelescope:2022urf; EventHorizonTelescope:2022xqj. Due to the strong gravitational lensing effect of the black hole, the surrounding light rays are captured and form the central shadow in the image Virbhadra:1999nm; Virbhadra:2002ju; Virbhadra:2007kw; Synge:1966okc; Bardeen:1972fi; Bozza:2010xqn. The boundary of the black hole shadow is determined by the critical impact parameter corresponding to the unstable circular orbit of the photon Perlick:2021aok. Moreover, the bright ring-like structure results from the superposition of multiple light rays. This is closely related to the distribution and structure of accretion matter, and strong gravitational lensing effects around the black hole.
@@
In fact, originally, the black hole shadow was termed the escape cone of the photon Synge:1966okc. Subsequently, the angular radius of the photon capture region of the Schwarzschild black hole was derived Luminet:1979nyg. Bardeen’s research further showed that a rotating Kerr black hole would produce a $D$-shaped shadow structure, which is attributed to the dragging effect of the black hole, resulting in an asymmetry in the orbital radii of prograde and retrograde photons Bardeen:1972fi. This has given rise to extensive discussion on black hole shadows, and various numerical simulations of black hole shadows have been discussed Shen:2005cw; Yumoto:2012kz; Atamurotov:2013sca; Papnoi:2014aaa; Abdujabbarov:2015xqa; Kumar:2018ple, which have shown that black hole shadows depend on spacetime geometry. In addition, black hole shadows in various modified theories of gravity and different spacetime dimensions have been extensively investigated, such as the Gauss-Bonnet theory Ma:2019ybz; Guo:2020zmf, conformal gravity Meng:2022kjs, the Chern-Simons type theory Meng:2023wgi; Ayzenberg:2018jip; Amarilla:2010zq, $f(R)$ gravity Addazi:2021pty; Dastan:2016vhb, and so on Amarilla:2011fx; Amarilla:2013sj; Amir:2017slq; Mizuno:2018lxz; Eiroa:2017uuq; Vagnozzi:2019apd; Banerjee:2019nnj; Chowdhuri:2020ipb; Konoplya:2019goy; Younsi:2016azx; Olmo:2023lil. Furthermore, the shadows of naked singularities Shaikh:2018lcc; Joshi:2020tlq; Dey:2020bgo and wormholes Rahaman:2021web; Kasuya:2021cpk; Shaikh:2018oul; Wielgus:2020uqz; Peng:2021osd; Neto:2022pmu; Tsukamoto:2012xs have been discussed, respectively. It is worth noting that the EHT has provided observational constraints on the shadow characteristics of the supermassive black holes M87* and SgrA*. For example, EHT observations have constrained the angular diameter of the SgrA* black hole shadow to be $\theta_{d}=48.7\pm 7\mu as$ EventHorizonTelescope:2022wkp, and the angular diameter of the supermassive M87* black hole shadow to be $29.32\mu as<\theta_{d}<51.06\mu as$ Kuang:2022ojj; EventHorizonTelescope:2021dqv. Although the supermassive black holes M87* and SgrA* are generally considered to be Kerr black holes, the limited resolution of the EHT still leaves room for alternative black hole models, or those predicted by modified theories of gravity. In addition, the astronomical observations of the supermassive black holes M87* and SgrA* can be used to test and constrain both alternative general relativistic black hole models and modified theories of gravity. Consequently, using the black hole shadow observed by the EHT to test and constrain the modified gravity theories has been widely discussed in Vagnozzi:2022moj; Afrin:2021imp; Kuang:2022ojj; Tang:2022hsu; Tang:2022bcm; Kuang:2022xjp; Kumar:2019pjp; Shaikh:2021yux; Wu:2023yhp; Capozziello:2023tbo; Sui:2023rfh; Pantig:2022qak; Ghosh:2023kge; Tsukamoto:2014tja.
@@
The bright ring-like structure in EHT images is closely associated with accreting matter around the black hole, and its spatial distribution is significantly influenced by the surrounding luminous accretion flow. This accretion material is mainly composed of gas and dust, forming a rapidly rotating accretion disk around the black hole. The formation of black hole images arises from the combined effects of accretion disk radiation, photon rings, gravitational lensing, and relativistic effects, which generally require numerical simulations using general relativistic magnetohydrodynamics (GRMHD) EventHorizonTelescope:2019pcy. In fact, some simple accretion disk models are sufficient to capture key features of black hole images and reveal the properties of strong gravitational field regions. Wald et al. first investigated the imaging features of Schwarzschild black hole using a simplified thin accretion disk model Gralla:2019xty, and defined three types of emission from the accretion disk based on the number $m$ of intersections of light rays with the accretion disk: direct emission ($m=1$), lensed ring emission ($m=2$), and photon ring emission ($m\geq 3$). Their study revealed that the observed intensity of the accretion disk primarily originates from direct emission. The contribution of lensed ring emission is relatively small, and the contribution of photon ring emission is negligible. Subsequently, the thin accretion disk models and spherically symmetric accretion models for various types of black holes have been widely studied Dokuchaev:2019pcx; Peng:2020wun; He:2021htq; Eichhorn:2021iwq; Li:2021riw; Wang:2023vcv; Wang:2024lte; Meng:2024puu; Zeng:2020dco; Capozziello:2025wwl; Alloqulov:2025pex; Alloqulov:2024zln; Saurabh:2020zqg; Zeng:2020vsj; Qin:2020xzu; Narayan:2019imo. These studies demonstrate that the black hole images are primarily determined by the properties of spacetime geometry rather than the details of accretion. It is noteworthy that black holes featuring a double photon sphere structure exhibit additional ring features in their images, which differs from the typical case of a single photon sphere Gan:2021xdl; Gan:2021pwu; Meng:2023htc; Wang:2025hzu. Moreover, considering more realistic astrophysical conditions, the accretion images of various types of rotating black holes have also been extensively studied Hou:2022eev; Wang:2023fge; Liu:2021yev; Zhang:2023bzv; He:2024amh; Zheng:2024ftk; Guo:2024mij; Heydari-Fard:2023kgf; Donmez:2024lfi; Yang:2024nin; Li:2024ctu; Wang:2024uda; Meng:2025ivb. In general, using black hole image features to distinguish between general relativistic black holes Guo:2022iiy; Guo:2022ghl; Archer-Smith:2020hqq; Okyay:2021nnh; Uniyal:2022vdu; Hou:2022eev; Uniyal:2023inx; Akbarieh:2023kjv; Gao:2023mjb; Theodosopoulos:2023ice; Meng:2023uws, other compact objects Boshkayev:2022vlv; Xavier:2023exm; Sakai:2014pga; Bacchini:2021fig; Destounis:2023khj, and black holes in modified gravity theories is a powerful detection method, but this does not always work, as even in GR, the images of black holes may exhibit degeneracies. For example, sufficiently dense boson stars Liebling:2012fv; Rosa:2023qcv may have shadows similar to those of classical black holes Rosa:2022tfv and images of very relativistic rotating boson stars may resemble Kerr black holes Vincent:2015xta.
@@
This paper is organized as follows. In Sec.II, we study the effective potential of photons around the black hole to analyze the photon sphere radius and the impact parameter, and also constrain the range of the coupling parameter by means of EHT observations. In Sec.III, utilizing the Euler-Lagrange equation and the ray-tracing method, we calculated the deflection trajectory of photons around the Bardeen black hole surrounded by PFDM. Furthermore, we employed two simplified emission functions to investigate the total observed intensity of black hole illuminated by the thin disk accretion. In Sec.IV, we presented the shadow and image of a black hole illuminated by a static spherical accretion. Sec.V presents the conclusion.
@@ II Photons sphere and observational constraints of Bardeen black hole surrounded by PFDM
For the static spherically symmetric spacetime, considering the coupling between the gravitational field and the non-linear electromagnetic field, a Bardeen black hole surrounded by PFDM can be obtained Zhang:2020mxi
@@
where $g$, $M$, and $\alpha$ are the magnetic charge, mass, and dark matter parameter of the black hole, respectively. In addition, to ensure the validity of the weak energy condition, the dark matter parameter $\alpha$ must satisfy $\alpha<0$ Sun:2023woa. When the dark matter parameter $\alpha=0$, the black hole reduces to the regular Bardeen black hole. For magnetic charge $g=0$, the black hole reduces to the Schwarzschild black hole surrounded by PFDM Li:2012zx; Ayon-Beato:2000mjt, and when $\alpha=0$ and $g=0$, the black hole reduces to the Schwarzschild black hole. In the following study, we will rescale all the physical quantities by the mass $M$ and set $M=1$ for simplicity. By solving equation $f(r)=0$, we can determine the event horizon $r_{h}$ of the black hole. However, since the explicit expressions for the solutions of the event horizon $r_{h}$ is too complicated, we show the relationship between the event horizon $r_{h}$ and the magnetic charge parameter $g$ and the dark matter parameter $\alpha$ in Fig.1. Obviously, the dark matter parameter $\alpha$ and the magnetic charge parameter $g$ exhibit distinctly different influences on the event horizon. As the coupling parameter $\left|\alpha\right|$ decreases or $g$ increases, the event horizon $r_{h}$ shows a monotonically decreasing trend.
Next, we investigate the null geodesic motion and shadow of the Bardeen black hole surrounded by PFDM. The trajectory of photon in curved spacetime can be described by the Euler-Lagrange equation as
@@
$E\equiv-\frac{\partial\mathcal{L}}{\partial\dot{t}}=f(r)\dot{t},~~~~~~~~~~L_{z}=\frac{\partial\mathcal{L}}{\partial\mathcal{\dot{\phi}}}=r^{2}\sin^{2}\theta\dot{\phi}.$ (4)
In addition, we consider the motions of photons moving on the equatorial plane $\theta=\pi/2$ without losing generality. Furthermore, considering $\mathcal{L}=0$ for the photon and defining the impact parameter $b=L_{z}/E$, we can obtain three first-order differential equations of motion for the photons
@@
where $b_{ph}$ is the critical impact parameter corresponding to the radius $r_{ph}$ of photon sphere. Subsequently, the relationship between the photon sphere radius $r_{ph}$ and the parameters (the magnetic charge $g$ and the dark matter parameter $\alpha$) is shown below in Fig.1. Similar to the event horizon, the smaller the dark matter parameter $\left|\alpha\right|$ or the larger magnetic charge parameter $g$, the smaller the corresponding photon sphere radius $r_{ph}$. In addition, we give the curves of critical impact parameters and effective potential changing with parameters $\alpha$ and $g$ in Fig.2. Obviously, a smaller $\left|\alpha\right|$ and a larger $g$ correspond to a smaller critical impact parameter $b_{ph}$, a larger peak in the effective potential, and a leftward shift of that peak. Overall, for the event horizon $r_{h}$, the radius of the photon sphere $r_{ph}$, and the effective potential $V_{\text{eff}}$, there is a competitive effect between the dark matter parameter $\left|\alpha\right|$ and the magnetic charge parameter $g$.
Figure 1: The event horizon $r_{h}$ and the photon sphere radius $r_{ph}$ as functions of the dark matter parameter $\alpha$ and the magnetic charge parameter $g$ of the Bardeen black hole surrounded by PFDM. Here we set $M=1$.
@@
On the other hand, for the distant observer, the angular diameter $\theta_{d}$ of black hole shadow can be given by Kumar:2020owy
@@
Figure 3: Left: the constraints of the coupling parameters $\alpha$ and $g$ by the observational shadow of M87* black hole. The red curve corresponds to the angular diameter $\theta_{d}=51.06~\mu as$. Right: the constraints of the coupling parameters $\alpha$ and $g$ by the observational shadow of SgrA* black hole. The red curve corresponds to the angular diameter $\theta_{d}=55.7~\mu as$.